# Extreme time extrapolation capabilities and thermodynamic consistency of physics-inspired Neural Networks for the 3D microstructure evolution of materials

### v1

@@ III.1 Domain size and evolution time generalization tests
One of the main advantages of a fully convolutional NN is the possibility of applying the trained model to inputs of arbitrary size. This is made possible thanks to the implicit assumption of locality embedded in the Convolutional layer choice. While for non-local PDEs this kind of procedure should be considered with care, the current application to the Cahn-Hilliard equation as defined in Eq. 2 does not pose any issues. To check the generalization capabilities of our CRNN in this setting, we report in Fig. 2(a) a computational domain twice as large in all spatial dimensions and evolved for double the amount of time steps used for training examples. It can be observed that the NN prediction closely reproduces the finite difference evolution at all reported stages.
Figure 3: Time generalization test. A domain enclosing a region in which $\varphi$ rapidly fluctuates is evolved for $40$ times more steps than those used in training. Despite local variations, long-time behavior still exhibits quantitative correspondence. In the upper half of the snapshots, only $\varphi\geq 0.5$ is shown.
As an additional proof of the CRNN generalization capabilities to qualitatively different configurations, in Fig. 2(b) we also report the evolution under the Cahn-Hilliard flow of a domain initialized with the shape of the Stanford bunny StanfordBunny , which is clearly out of the distribution of typical training set microstructures. The domain size, in this case, is the same as the $64\times 64\times 64$ grid used in training, but coarsening stages at 100 $\tau$ are inspected as in panel (a). Close correspondence with the ground truth evolution is again achieved, despite some local variations (e.g. the discrepancy at 20 $\tau$ in the left ear), which however are not preventing the close correspondence at later stages.
We remark that, similarly to a simple forward-Euler integration scheme, computational costs scale linearly with the number of collocation points, as explicitly reported in Appendix A, albeit with a longer time interval between subsequent predicted states.
@@
Another generalization test regards the possibility of generating sequences of arbitrary length by iteratively re-feeding the CRNN’s own output as input to generate the following evolution stage. This aspect is particularly critical, as this recurrence may lead to prediction error accumulation and loss of reliability in the generated sequence. In principle, therefore, one could expect the evolution obtained by NN and the one coming from finite difference integration to diverge gradually. Furthermore, as dynamics progress, predicted morphologies arising from domain coarsening might be less represented in the training set. This could lead to particularly severe limitations for ML approaches, as configurations emerging from the considered Phase Field model are subject to several constraints, in particular the conservation and boundedness of $\varphi$. Notice that, while local conservation is ensured (up to numerical precision) by the specialized, physics-inspired layer ("toFlux" and divergence operation in Fig. 1(b)), there is no guarantee, so far, that very long sequences will not produce invalid phase field representations, e.g. with diverging $\varphi$ values. To demonstrate the generalization capabilities of the proposed CRNN, in Fig. 3(a) we show the comparison between the predicted and ground truth evolution for an initial configuration with fluctuations localized in the central slice of a $128\times 128\times 128$ grid. It can be seen that there is a substantial agreement between such "long-time" predictions from the NN and the true evolutions provided by the explicit integration scheme. Once again, despite local differences, the overall sequence and the final state of the evolution exhibit compelling correspondence. Notice that the complete evolution requires $8\times 10^{5}$ finite difference integration steps, which correspond to a sequence $40$ times longer than those contained in the training set.
@@ III.2 Stationary state prediction
As stated in previous Sections, CRNN approaches are particularly convenient from a computational standpoint, as numerical operations can be easily parallelized on GPUs and multi-threaded systems. This allows for a strong compression of simulation wall times. Regarding examples in previous sections, the speedup obtained on an Nvidia RTX A4000 GPU running on the same machine that performed the (non-optimized) Phase Field simulations was of $\approx 10^{3}$. This is only an indicative figure of merit, as, for instance, more advanced integration schemes could significantly speed up spinodal decomposition simulations CHEN1998PFpseudospectral . On the other hand, a similar NN approach could be applied to more complex PDEs, where such advanced schemes are possibly not available or not as advantageous.
@@
We now take profit from this computational speedup to inspect what happens if the NN model is run to generate sequences of orders of magnitude longer than those provided in training. In particular, our scope is here to investigate whether the error accumulation eventually leads the NN predictions to un-physical configurations and whether the predicted stationary states (if any) are consistent with those expected from the Cahn-Hilliard model. Fig.4 reports the NN-predicted sequence of microstructure coarsening stages going from a Perlin-noise profile (with similar parameters to those used in training but not present in the dataset) to its stationary state, together with a one-to-one comparison with the corresponding finite difference integration up to $\approx 1/3$ of the evolution. Notably, the generated sequence requires $8\times 10^{4}$ NN iterations, which is 4 orders of magnitude longer than those provided in training. A full simulation using the explicit integration scheme therefore requires more computational resources than those used for the whole dataset building.
@@
Regarding error accumulation, local discrepancies reduce as dynamics progress, as expected from the Cahn-Hilliard flow minimizing the interface area between the two phases. Inspecting the configurations at time $1.2\times 10^{4}\tau$ of Fig. 4, it is possible to notice that the connectivity of the red domain shows significant deviations (see also close-ups provided in the insets at the bottom of the figure). At the time $2.4\times 10^{4}\tau$, however, the difference between the two configurations is far less severe, showing that the NN prediction has closed the gap with the ground-truth evolution. This testifies that the CRNN has correctly learned the underlying PF model and that no net error accumulation is pushing the predictions systematically toward un-physical configurations. This can be, at least in part, ascribed to the Recurrent nature of the NN: since the model is trained to minimize the likelihood of a whole sequence, temporary discrepancies are allowed, as long as the overall prediction gets closer again to the ground truth in the long run. A quantification of this effect is provided in Sect. III.3.
@@
Fig. 4 also reports a stationary state presenting a layered microstructure, which is reached at time $\approx 8\times 10^{4}\tau$ and corresponds to a time sequence $1600$ times longer than those observed in training. Importantly, such configuration is stable once reached, despite being out of the distribution of the examples in the dataset. This is a striking proof of the generalization capabilities of the CRNN, as this layered microstructure is one of the expected stationary solutions of the Cahn-Hilliard equation under the Periodic Boundary Conditions considered here provatas2011phase .
@@ III.3 Quantitative evaluation of predictive performances
Until this point, the assessment of the NN performances was only qualitative and based on visual inspection. This last section addresses this issue by performing a more quantitative analysis. The long-time behavior of the system is investigated using a spinodal decomposition simulation on a $128\times 128\times 128$ grid, starting from a Perlin noise initial condition and $7.2\times 10^{4}\tau$ long, compatible with the one already shown in Fig. 4. The comparison between representative stages of the two sequences is reported in Fig. 5(a). To quantify the difference between the two evolutions, we report the Root Mean Squared Error (RMSE), here evaluated as
@@
Despite the initial condition being the same for both the NN and the ground truth evolution, the RMSE starts from a non-vanishing value. The origin of such behavior is due to the artifacts introduced by the downscaling interpolation required by the NN. The initial RMSE of $\approx 0.13$ should therefore be considered "intrinsic" to the downscaling procedure, but it does not affect the subsequent evolution. This is verified by calculating the same quantity after interpolation on a $64\times 64\times 64$ grid, which is reported in Appendix B. In later evolution stages, the RMSE increases and decreases again, e.g. at $\tau\approx 1.8\times 10^{4}$ and at the end of the reported evolution, when both the NN prediction and the ground truth evolution reach a layered microstructure. This shows that the accumulation of deviations is progressively recovered, as expected from the Cahn-Hilliard evolution. It is also interesting to note that the time required to reach a two-interface microstructure is consistent between the true and predicted evolution, although the CRNN is faster in smoothing interface undulations.
@@
If the ground truth $\varphi$ and the prediction $\hat{\varphi}$ differ by a small displacement of the position of the interface between the two phases, Eq. 4 yields the root mean square value of such displacement. A more complete discussion may be found in Appendix C. $\bar{\delta}$ can therefore be interpreted as the average, local distance between the predicted and true interface in grid units. Similarly to what was observed for the RMSE, this value increases in the intermediate stages and decreases towards the end of the reported dynamics, as the microstructure coarsens forming interfaces parallel to the computational domain sides. The initial condition high value in the RMSE is far less pronounced in terms of $\bar{\delta}$, but notice that this value refers to a random noise configuration lacking well-defined interfaces. We recall that the interface thickness parameter $\varepsilon=3$ (see Eq. 2), which means that the root mean square distance between the interfaces is comparable for most of the evolution to the width of transition regions between the two phases.
@@
A point-by-point comparison of the evolution yields quantitative information on the NN prediction on a fine level. Indeed, both the RMSE and $\bar{\delta}$ are sensitive to local variations. However, they do not provide information on how much the predicted dynamics are globally consistent with the underlying physical mechanisms. In this respect, it is crucial to assess whether the CRNN prediction remains consistent with the Cahn-Hilliard flow even when the RMSE or $\bar{\delta}$ values are at their maxima. Fig. 6(a) reports the time dependence of the free energy of the system normalized on the domain volume $F/N$ for both the NN prediction and the corresponding ground-truth evolution. A logarithmic scale is used to appreciate finer details and to analyze the power-law trends of free-energy decay. In panel (b), the relative absolute error is also reported (linear scale is used in this case). It is evident how the predicted evolution of the free energy closely reproduces the "true" curve, even when the error values of Fig. 5(a) increase, testifying that the CRNN predictions follow the underlying free-energy minimization process. The largest deviations are observed near the late stages of the evolution ($log(\tau)\geq 10.25$, i.e. $\tau\geq 2.8\times 10^{4}$), which exhibit an acceleration in the coarsening rate with respect to the true evolution. Nonetheless, the relative absolute error in the predicted $F$ stays below 5.5% for the whole evolution, as shown in Fig. 5(b).
@@
Fig. 6(a) also allows for a second quantitative analysis, regarding the coarsening rate of the microstructure. In Cahn-Hilliard evolution the typical domain length is expected to grow in time as $t^{1/3}$ and, as a consequence, the interface area per unit volume is expected to decay as $t^{-1/3}$ kohn2002upper ; kwon2010morphology . From the definition of $F$ in Eq. 1, in this work, the free energy corresponds to the interface area between the two phases. We therefore expect that $F(t)\propto t^{-1/3}$ in the long-time regime. The dashed black line reported in Fig. 6(a) corresponds to such behavior. Clearly, the ground truth $F(t)$ line is parallel to the expected one, indicating the same dependence for the free energy decay. Local deviations and fluctuations are related to finite domain size effects and interface splitting/merging processes. The NN prediction also follows the same power-law, up to $log(\tau)\approx 10.25$ (again, corresponding to $\tau\approx 2.8\times 10^{4}$), where the coarsening rate increases with respect to the ground truth and gets closer to a $\propto t^{-0.44}$ dependence (not shown). This behavior is probably due to the lack of long-time configurations in the training set and may be eliminated with active or transfer learning workflows, the implementation of which we leave for future works. Remarkably, while not completely satisfactory, this does not prevent the NN from predicting the correct stationary state.
@@
To confirm the statistical significance of the above discussion, we perform a similar analysis on an ensemble of 100 evolutions, obtained using a $64\times 64\times 64$ grid and Perlin-noise independent initial configurations consistent with those in the training set. A total time of $\tau=500$ is considered. Fig. 6(c) reports in semi-logarithmic scale the decay of the free energy obtained by averaging across all evolutions, along with its variability range, represented by the standard deviation and visualized as the shaded area. A close-up of the two curves is also shown to better appreciate fine details. As it can be seen, predicted and ground-truth quantities are in close agreement, both in terms of means and standard deviations, with the largest difference in the initial stages, where the effects of the random initialization are stronger. These results confirm that the generated microstructures are quantitatively consistent with the underlying Cahn-Hilliard material flow also on an ensemble level.
@@
We close the discussion by addressing the problem of the wrong exponent in the free energy dissipation observed at the end of the evolution in Fig. 6(a). While performing a CRNN-ground truth comparison is computationally demanding, we can attack the problem indirectly by exploiting the low cost of Machine Learning predictions. Fig. 6(d) reports in logarithmic scale the average free energy decay curve over 10 independent evolutions $80000\tau$ long on a $128\times 128\times 128$ grid, each starting from a different Perlin noise initial condition. One standard deviation interval is also reported with the red-shaded area. As it can be clearly appreciated, the average long-time behavior exhibits the expected $t^{-1/3}$ power law, showing that the acceleration previously reported is not systematic. It may therefore be concluded that the NN predictions do not exhibit a net, strong deviation of the expected behavior of $F(t)$, despite the relatively short sequences used in training.

### v2

@@ III.1 Domain size and evolution time generalization tests
One of the main advantages of a fully convolutional NN is the possibility of applying the trained model to inputs of arbitrary size. This is made possible thanks to the implicit assumption of locality embedded in the Convolutional layer choice. While for non-local PDEs this kind of procedure should be considered with care, the current application to the Cahn-Hilliard equation as defined in Eq. 2 does not pose any issues. To check the generalization capabilities of our CRNN in this setting, we report in Fig. 2(a) a computational domain twice as large in all spatial dimensions and evolved for double the amount of time steps used for training examples. It can be observed that the NN prediction closely reproduces the finite difference evolution at all reported stages. The MSE loss on the whole sequence is also reported, exhibiting a value slightly higher than training and validation.
Figure 3: Time generalization test. A domain enclosing a region in which $\varphi$ rapidly fluctuates is evolved for $40$ times more steps than those used in training. Despite local variations, long-time behavior still exhibits quantitative correspondence. In the upper half of the snapshots, only $\varphi\geq 0.5$ is shown. MSE loss for the whole sequence is reported in the inset.
As an additional proof of the CRNN generalization capabilities on qualitatively different configurations, in Fig. 2(b) we also report the evolution under the Cahn-Hilliard flow of a domain initialized with the shape of the Stanford bunny Sta (2023), which is clearly out of the distribution of typical training set microstructures. The domain size, in this case, is the same as the $64\times 64\times 64$ grid used in training, but coarsening stages at 100 $\tau$ are inspected as in panel (a). Close correspondence with the ground truth evolution is again achieved, despite some local variations (e.g. the discrepancy at 20 $\tau$ in the left ear), which however are not preventing the close correspondence at later stages. The MSE loss, reported at the bottom of Fig. 2 is comparable to values in training/validation, confirming the excellent generalization capabilities of the CRNN.
We remark that, similarly to a simple forward-Euler integration scheme, computational costs scale linearly with the number of collocation points, albeit with a longer time interval between subsequent predicted states. We point the technical reader to Appendix A where this aspect is quantitatively analyzed.
@@
Another generalization test regards the possibility of generating sequences of arbitrary length by iteratively re-feeding the CRNN’s own output as input to generate the following evolution stage. This aspect is particularly critical, as this recurrence may lead to prediction error accumulation and loss of reliability in the generated sequence. In principle, therefore, one could expect the evolution obtained by NN and the one coming from finite difference integration to diverge gradually. Furthermore, as dynamics progress, predicted morphologies arising from domain coarsening might be less represented in the training set. This could lead to particularly severe limitations for ML approaches, as configurations emerging from the considered Phase Field model are subject to several constraints, in particular the conservation and boundedness of $\varphi$. Notice that, while local conservation is ensured (up to numerical precision) by the specialized, physics-inspired layer ("toFlux" and divergence operation in Fig. 1(b)), there is no guarantee, so far, that very long sequences will not produce invalid phase field representations, e.g. with diverging $\varphi$ values. To demonstrate the generalization capabilities of the proposed CRNN, in Fig. 3(a) we show the comparison between the predicted and ground truth evolution for an initial configuration with fluctuations localized in the central slice of a $128\times 128\times 128$ grid. It can be seen that there is a substantial agreement between such "long-time" predictions from the NN and the true evolutions provided by the explicit integration scheme. Once again, despite local differences, the overall sequence and the final state of the evolution exhibit compelling correspondence, with an MSE loss of the same order of magnitude as the one observed in the top evolution of Fig. 2. Notice that the complete evolution requires $8\times 10^{5}$ finite difference integration steps, which correspond to a sequence $40$ times longer than those contained in the training set.
@@ III.2 Stationary state prediction
As stated in previous Sections, CRNN approaches are particularly convenient from a computational standpoint, as numerical operations can be easily parallelized on GPUs and multi-threaded systems. This allows for a strong compression of simulation wall times. Regarding examples in previous sections, the speedup obtained on an Nvidia RTX A4000 GPU running on the same machine that performed the (non-optimized) Phase Field simulations was of $\approx 10^{3}$. This is only an indicative figure of merit, as, for instance, more advanced integration schemes could significantly speed up spinodal decomposition simulations Chen and Shen (1998). On the other hand, a similar NN approach could be applied to more complex PDEs, where such advanced schemes are possibly not available or not as advantageous.
@@
We now take profit from this computational speedup to inspect what happens if the NN model is run to generate sequences orders of magnitude longer than those provided in training. In particular, our scope is here to investigate whether the error accumulation eventually leads the NN predictions to un-physical configurations and whether the predicted stationary states (if any) are consistent with those expected from the Cahn-Hilliard model. Fig.4 reports the NN-predicted sequence of microstructure coarsening stages going from a Perlin-noise profile (with similar parameters to those used in training but not present in the dataset) to its stationary state, together with a one-to-one comparison with the corresponding finite difference integration up to $\approx 1/3$ of the evolution. Notably, the generated sequence requires $8\times 10^{4}$ NN iterations, which is 4 orders of magnitude longer than those provided in training. A full simulation using the explicit integration scheme therefore requires more computational resources than those used for the whole dataset building.
@@
Regarding error accumulation, local discrepancies reduce as dynamics progress, as expected from the Cahn-Hilliard flow minimizing the interface area between the two phases. Indeed, small dimples or corrugations in a lower curvature interface tend to be flattened by the evolution law defined by Eq. 2, as can be observed in the initial stages of the evolution (see e.g. fine details in the Stanford bunny in Fig. 2). Inspecting the configurations at time $1.2\times 10^{4}\tau$ of Fig. 4, it is possible to notice that the connectivity of the red domain shows significant deviations (see also close-ups provided in the insets at the bottom of the figure). At the time $2.4\times 10^{4}\tau$, however, the difference between the two configurations is far less severe, showing that the NN prediction has closed the gap with the ground-truth evolution. This testifies that the CRNN has correctly learned the underlying PF model and that no net error accumulation is pushing the predictions systematically toward un-physical configurations. This can be, at least in part, ascribed to the Recurrent nature of the NN: since the model is trained to maximize the likelihood of a whole sequence, temporary discrepancies are allowed, as long as the overall prediction gets closer again to the ground truth in the long run. A quantification of this effect is provided in Sect. III.3.
@@
Fig. 4 also reports a stationary state presenting a layered microstructure, which is reached at time $\approx 8\times 10^{4}\tau$ and corresponds to a time sequence $1600$ times longer than those observed in training. Importantly, such configuration is stable once reached, despite being out of the distribution of the examples in the dataset. This is a striking proof of the generalization capabilities of the CRNN, as this layered microstructure is one of the expected stationary solutions of the Cahn-Hilliard equation under the Periodic Boundary Conditions considered here Provatas and Elder (2011).
@@ III.3 Quantitative evaluation of predictive performances
Until this point, the assessment of the NN performances was only qualitative and mainly based on visual inspection. This last section addresses this issue by performing a more quantitative analysis. The long-time behavior of the system is investigated using a spinodal decomposition simulation on a $128\times 128\times 128$ grid, starting from a Perlin noise initial condition and $7.2\times 10^{4}\tau$ long, compatible with the one already shown in Fig. 4. The comparison between representative stages of the two sequences is reported in Fig. 5(a). To quantify the difference between the two evolutions, we report the Root Mean Squared Error (RMSE), here evaluated as ($V$ is the computational domain volume)
@@
Notice that, despite the initial condition being the same for both the NN and the ground truth evolution, the RMSE starts from a non-vanishing value. The origin of such behavior is due to the artifacts introduced by the downscaling interpolation required by the NN. The initial RMSE of $\approx 0.13$ should therefore be considered "intrinsic" to the downscaling procedure, but it does not affect the subsequent evolution. In Appendix B, we explicitly verify this by calculating the same quantity after interpolation on a $64\times 64\times 64$ grid. In later evolution stages, the RMSE increases and decreases again, e.g. at $\tau\approx 1.8\times 10^{4}$ and at the end of the reported evolution, when both the NN prediction and the ground truth evolution reach a layered microstructure. This shows that the accumulation of deviations is progressively recovered, as expected from the Cahn-Hilliard evolution. It is also interesting to note that the time required to reach a two-interface microstructure is consistent between the true and predicted evolution, although the CRNN is faster in smoothing interface undulations.
@@
If the ground truth $\varphi$ and the prediction $\hat{\varphi}$ differ by a small displacement of the position of the interface between the two phases, Eq. 4 yields the root mean square value of such displacement. Appendix C reports a more complete discussion and a mathematical derivation. $\bar{\delta}$ can therefore be interpreted as the average, local distance between the predicted and true interface in grid units. Similarly to what was observed for the RMSE, this value increases in the intermediate stages and decreases towards the end of the reported dynamics, as the microstructure coarsens forming interfaces parallel to the computational domain sides. The initial condition high value in the RMSE is far less pronounced in terms of $\bar{\delta}$, but notice that this value refers to a random noise configuration lacking well-defined interfaces. We recall that the interface thickness parameter $\varepsilon=3$ (see Eq. 2), which means that the root mean square distance between the interfaces is comparable for most of the evolution to the width of transition regions between the two phases.
@@
A point-by-point comparison of the evolution yields quantitative information on the NN prediction on a fine level. Indeed, both the RMSE and $\bar{\delta}$ are sensitive to local variations. However, they do not provide information on how much the predicted sequences are globally consistent with the underlying physical mechanisms. In this respect, it is crucial to assess whether the CRNN prediction remains consistent with the Cahn-Hilliard flow even when the RMSE or $\bar{\delta}$ values are at their maxima. Fig. 6(a) reports the time dependence of the free energy of the system normalized on the domain volume $F/N$ for both the NN prediction and the corresponding ground-truth evolution. A logarithmic scale is used to appreciate finer details and to analyze the power-law trends of free-energy decay. In panel (b), the relative absolute error is also reported (linear scale is used in this case). It is evident how the predicted evolution of the free energy closely reproduces the "true" curve, even when the error values of Fig. 5(a) increase, testifying that the CRNN predictions follow the underlying free-energy minimization process. The largest deviations are observed near the late stages of the evolution ($log(\tau)\geq 10.25$, i.e. $\tau\geq 2.8\times 10^{4}$), which exhibit an acceleration in the coarsening rate with respect to the true evolution. Nonetheless, the relative absolute error in the predicted $F$ stays below 5.5% for the whole evolution, as shown in Fig. 5(b).
@@
Fig. 6(a) also allows for a second quantitative analysis, regarding the coarsening rate of the microstructure. In Cahn-Hilliard evolution the typical domain length is expected to grow in time as $t^{1/3}$ and, as a consequence, the interface area per unit volume is expected to decay as $t^{-1/3}$ Kohn and Otto (2002); Kwon et al. (2010). From the definition of $F$ in Eq. 1, in this work, the free energy corresponds to the interface area between the two phases. We therefore expect that $F(t)\propto t^{-1/3}$ in the long-time regime. The dashed black line reported in Fig. 6(a) corresponds to such behavior. Clearly, the ground truth $F(t)$ line is parallel to the expected one, indicating the same dependence for the free energy decay. Local deviations and fluctuations are related to finite domain size effects and interface splitting/merging processes. The NN prediction also follows the same power-law, up to $log(\tau)\approx 10.25$ (again, corresponding to $\tau\approx 2.8\times 10^{4}$), where the coarsening rate increases with respect to the ground truth and gets closer to a $\propto t^{-0.44}$ dependence (not shown). This behavior is probably due to the lack of long-time configurations in the training set and may be eliminated with active or transfer learning workflows, the implementation of which we leave for future works. Remarkably, while not completely satisfactory, this does not prevent the NN from predicting the correct stationary state.
@@
To confirm the statistical significance of the above discussion, we perform a similar analysis on an ensemble of 100 evolutions, obtained using a $64\times 64\times 64$ grid and Perlin-noise independent initial configurations consistent with those in the training set. A total time of $\tau=500$ is considered, as the computational effort required makes it unpractical to consider longer sequences. Fig. 6(c) reports in semi-logarithmic scale the decay of the free energy obtained by averaging across all evolutions, along with its variability range, represented by the standard deviation and shown as the shaded area. A close-up of the two curves is also shown to better appreciate fine details. As it can be seen, predicted and ground-truth quantities are in close agreement, both in terms of means and standard deviations, with the largest difference in the initial stages, where the effects of the random initialization are stronger. These results confirm that the generated microstructures are quantitatively consistent with the underlying Cahn-Hilliard material flow also on an ensemble level.
@@
We close the discussion by addressing the problem of the wrong exponent in the free energy dissipation observed at the end of the evolution in Fig. 6(a). While performing an extensive CRNN-ground truth comparison is computationally demanding, we can attack the problem indirectly by exploiting the low cost of Machine Learning predictions. Fig. 6(d) reports in logarithmic scale the average free energy decay curve over 10 independent evolutions $80000\tau$ long on a $128\times 128\times 128$ grid, each starting from a different Perlin noise initial condition. One standard deviation interval is also reported with the red-shaded area. As it can be clearly appreciated, the average long-time behavior exhibits the expected $t^{-1/3}$ power law, showing that the acceleration previously reported is not systematic. It may therefore be concluded that the NN predictions do not exhibit a net, strong deviation of the expected behavior of $F(t)$, despite the relatively short sequences used in training.